OpenStudy (anonymous):

how to solve x^2+14x=-24

OpenStudy (anonymous):

add a +24 on both sides of the equation to get \[x^2 +14x + 24 = 0\] then factor the equation to get \[(x + 12)(x+2) = 0\] you should be able to do the rest to figure out what x is...if you still need assistance, set \[x + 12 = 0\] and \[x + 2 = 0\] you should now be able to determine what x is